OpenStudy (anonymous):

: 3(x − 1) − (x + 5) = 2

OpenStudy (theviper):

\[3x-3-x-5=2\]

OpenStudy (theviper):

\[2x-8=2\]

OpenStudy (anonymous):

3(x − 1) − (x + 5) = 2 3x - 3 - x -5 = 2 2x = 10 x=5

OpenStudy (theviper):

\[2x=2+8=10\]

OpenStudy (anonymous):

Whoops, sorry viper, didn't notice you writing it out :/

OpenStudy (theviper):

\[\LARGE{\color{red}{x=\frac{10}{2}=5}}\]

OpenStudy (theviper):

NP:)